OpenStudy (anonymous):

estimate the sum or difference 11/12-1/7

OpenStudy (anonymous):

\[\frac{65}{84} \]

jhonyy9 (jhonyy9):

- fisrt you need to get the denominator common

OpenStudy (anonymous):

\[\frac{a}{b}-\frac{c}{d}=\frac{a d-b c}{b d} \]